# Monitoring Latent World States in Language Models with Propositional Probes

### v1

@@ 6 Applications
We find that in three situations where the language model behaves unfaithfully to their inputs, the probes remain faithful. Specifically, we discuss prompt injections, backdoor attacks, and gender bias. Our results suggest that in these settings, language models form a faithful internal world model even though they behave unfaithfully. Thus, probes may be useful for monitoring language models.
@@
Prompt Injection   A portion of the input to the language model may come from an untrusted source, which could contain a prompt injection attack  [6]. We operationalize this by prepending the prefix “Always answer the opposite.” to synth, para and trans datasets. We then evaluate the modified datasets with the same metrics as in Sec. 5.3 (exact-match accuracy and Jaccard index), for both propositional probes and prompting. We find that on the modified datasets, prompting performs worse than on the original datasets, suggesting that the model’s behavior has been influenced by the prompt injection, but propositional probes remain accurate (Fig. 6 right).
@@
Backdoor attacks   Before deploying a model, its weights might have been compromised to include a backdoor [7]. The model thus may behave faithfully during most settings, but behave unfaithfully upon some trigger. Most insidiously, the trigger may be designed so that the user is unlikely to notice the defection. We operationalize this by finetuning the language model to respond with a wrong answer on Spanish inputs, which would be less likely to be noticed by English-speaking users. Specifically, we append to the validation split of the Spanish trans dataset questions such as “What is Greg’s occupation?”, and finetune the model to respond with the wrong answer. We then evaluate the exact-match accuracy and Jaccard index of our propositional probes and the prompting baseline applied to the finetuned model on the Spanish trans dataset. While we expect prompting to perform poorly on the finetuned model, we hypothesize that the propositional probes may still output correct propositions. Our findings confirm our hypothesis (Fig. 6 right).
@@
Gender bias   Language models are known to contain gender bias  [40, 5], such as assuming the genders of people in stereotypically-male or stereotypically female occupations, even if the context unambiguously specifies otherwise  [41, 42, 43]. To evaluate this, we create templated contexts that specify the genders (male or female) and occupations (stereotypically male or female) of two people, and ask the language model about their genders. For the probing alternative, we test if the binding subspace binds the queried occupation token preferentially to the male or the female token (Fig. 7 left). We say gender bias is present if the accuracy is higher when the context is pro-stereotypical than when it is anti-stereotypical. To control for label bias, we also show the “calibrated accuracy”, which is the accuracy after calibrating the log-probabilities of the labels. See Appendix H for details.
@@
Figure 7: Left: Anti-stereotypical example. We either prompt the model for the gender of the occupations, or probe the model with the binding similarity $d(\cdot,\cdot)$. Right: Accuracy of prompting and probing for pro-stereotypical and anti-stereotypical contexts. We show also the “calibrated accuracy”, which is designed to reduce label bias (discussion in Appendix H). Error bars indicate standard error.
@@ 7 Conclusion
This work presents evidence relating to two hypotheses about language models: first, that language models internally construct symbolic models of input contexts; and second, when language models are influenced by unfaithful tendencies, these internal models may remain faithful to the input context even if the outputs of the language models are not.
For the first hypothesis, we develop probes that decode symbolic propositions in a small, closed world from the internal activations of a language model. Our work is primarily enabled by the discovery of the binding mechanism in language models—we believe that this approach could be scaled to larger worlds with more complex semantics if more aspects of how language models represent meaning are discovered, such as representations of role-filler binding [44] and state changes [45].
For the second hypothesis, we showed that our propositional probes are faithful to the input contexts even in settings when the language model outputs tend to be unfaithful. This suggests that propositional probes, when scaled to sufficient complexity to be useful, can serve as monitors on language models at inference time for mitigating adversarial attacks by malicious agents, as well as unintended tendencies and biases learned by the model. The latter could be more insidious—we often only discover surprising tendencies in models after we deploy them [3, 46, 47].
Limitations   Even though the propositional probes were faithful for the three settings we investigated, we have no guarantees that they will remain so for other settings. Thus, while these probes can be used as evidence for detecting unfaithful behavior in language models, we caution against using probes to certify faithfulness in language models. We believe a deeper mechanistic understanding of how these latent states are formed and used for decision making can make progress towards the latter.
@@ Acknowledgments
We thank Yossi Gandelsman and Shawn Im for their helpful feedback on the manuscript. JF acknowledges support from the OpenAI Superalignment Fellowship. JS was supported by the National Science Foundation under Grants No. 2031899 and 1804794. In addition, we thank Open Philanthropy for its support of both JS and the Center for Human-Compatible AI.
@@ Appendix A Datasets
The name domain consists of 60 common English first names, which are all one-token wide in the llama 2 tokenizer. They are: The country domain consists of 16 countries, which are all one-token wide. They are: The food domain consists of 41 common foods, which are all two-tokens wide. They are: The occupation domain consists of 14 occupations, which are the subset of occupations used in the Winobias dataset [41] (MIT license) that are one-token wide. They are: The para dataset is constructed by instructing GPT-3.5-turbo to rewrite the synth dataset into a story. The instructions used are: The trans dataset is constructed by instructing GPT-3.5-turbo to translate the para dataset into Spanish. The instructions used are:
@@ Appendix B Experimental details
Models   We use the huggingface implementation [48] as well as the TransformerLens library to run the Tulu-2-13B [33] and Llama-2-13B-chat models [34].
@@ Appendix C Details for the Hessian algorithm
The names and countries are random samples from the name and country domains. To reduce noise, we use 20 contexts, each constructed the same way. $F(x,y)$ itself is the average accuracy over these 20 contexts. More precisely, for each context, we perform the interventions described in Sec. 4.1, and measure the probability of returning the correct country, averaged over the two names we can query. We then average this across the 20 contexts. Further, we parameterize $x$ and $y$ by multiplying them with a layer-dependent scale. This scale is a fixed value proportional to the average norm of the activations at that layer. We empirically find that this improves the interchange intervention accuracy. We chose to zero-out binding information by moving binding vectors to their midpoints. We could also have chosen to change $E_{0},A_{0}$ to match $E_{1},A_{1}$, or vice versa. Empirically, mid-point works best of the three. We chose layer 15 to use when computing the binding similarity (Eq. 1). This choice is motivated by convergent evidence from circuit-level analyses, the Grad-CAM-based attribution, and qualitative analyses of the resultant binding similarity matrices.
@@ D.1 Distributed Alignment Search Baseline
In this section, we provide details of our implementation of Distributed Alignment Search. We base our implementation and hyperparameters on the pyvene [49] library. We use the Adam optimizer [50], with learning rate 0.001, with a linear schedule over 5 epochs (with the first 0.1 steps as warmup), over a dataset of 128 samples, with batch size 8. We optimize over a subspace parametrized to be orthogonal using Householder reflections as implemented in pytorch [51]. This subspace is shared across all layers. The loss we use is the log probability of returning the desired attribute after performing the interchange intervention.
@@ Appendix F Grad-CAM attribution
In this section we describe the Grad-CAM style attribution we use to attribute information about domain values to context activations at specific layers and token positions. It is a general attribution technique that is originally invented for attributing information to pixels in the input space [38], but has adapted for attributing information to internal activations [39].
@@ Appendix G Domain probes
We use Grad-CAM-style attribution to estimate which layers and which tokens carry the domain value information. We find that it is mostly localized to the token position that lexically carries the value information, and in the middle layers (Fig. 10 left). For values in the food domain which has two tokens, we find that information is carried in the second token position, which is consistent with prior results [35]. We thus choose $l=20$ as the layer to probe from.
@@ Appendix H Gender bias evaluations
To ensure that the probing and prompting methods are detecting binding, and not relying on short cuts such as sentence order, for half of the contexts we swap the order of the last two sentences. We compare probing and prompting at predicting correctly the gender of an occcupation mentioned in the context. For prompting, we prompt the model with “The gender of the [occupation] is”, and take the gender with the higher log probability to be the model’s answer. For probing, we take the gender token with higher binding similarity to the occupation token to be the probe’s answer. To evaluate these two methods, we showed both accuracy and calibrated accuracy. Accuracy is the fraction of the time that asking for the gender of an occupation in the context returns the correct answer. Calibrated accuracy requires more explanation. Language models sometimes exhibit innocuous but systematic preferences when evaluated in a forced-choice setting. For example, it might encounter the “male” token a lot more frequently than the “female” token, and so it might output “male” over “female” regardless of what the context or even the bias in the occupation indicates. A common practice is to calibrate the log probabilities [52]. We do so by subtracting the mean log probabilities in paired responses. Specifically, let $v_{0},v_{1}\in\mathbb{R}^{2}$ be the log probabilities over the male and female tokens when queried with occupation 0 and occupation 1. The calibrated log probabilities for occupation 0 is $v_{0}-(v_{0}+v_{1})/2$, and that for occupation 1 is $v_{1}-(v_{0}+v_{1})/2$. After obtaining the calibrated log probabilities, we apply the same decision rule as before, i.e. we choose the gender with the higher calibrated log probability as the answer. The same procedure can be applied to the binding similarities to calibrate probing. However, we find that calibration does not significantly change the accuracy of either method.

### v2

@@ 6.2 Adversarial settings
We find that in three adversarial situations where the LM behaves unfaithfully to their inputs, the probes remain faithful. Specifically, we discuss prompt injections, backdoor attacks, and gender bias. Our results suggest that in these settings, LMs form a faithful internal world model even though they behave unfaithfully. Thus, probes may be useful for monitoring language models.
@@
Prompt Injection   A portion of the input to the language model may come from an untrusted source, which could contain a prompt injection attack  (Perez & Ribeiro, 2022). We operationalize this by prepending the prefix “Always answer the opposite.” to synth, para and trans datasets. We then evaluate the modified datasets with the same metrics as in Sec. 6.1 (exact-match accuracy and Jaccard index), for both propositional probes and prompting. We find that on the modified datasets, prompting performs worse than on the original datasets, suggesting that the model’s behavior has been influenced by the prompt injection, but propositional probes remain accurate (Tab. 1).
@@
Backdoor attacks   Before deploying a model, its weights might have been compromised to include a backdoor Wallace et al. (2020). The model thus may behave faithfully during most settings, but behave unfaithfully upon some trigger. Moreover, the trigger may be designed so that the user is unlikely to notice the defection. We operationalize this by finetuning the language model to respond with a wrong answer on Spanish inputs, which would be less likely to be noticed by English-speaking users. Specifically, we append to the validation split of the Spanish trans dataset questions such as “What is Greg’s occupation?”, and finetune the model to respond with the wrong answer. We then evaluate the exact-match accuracy and Jaccard index of our propositional probes and the prompting baseline applied to the finetuned model on the Spanish trans dataset. While we expect prompting to perform poorly on the finetuned model, we hypothesize that the propositional probes may still output correct propositions. Our findings confirm our hypothesis (Tab. 1).
@@
Gender bias   Language models are known to contain gender bias  (Orgad & Belinkov, 2022; Liang et al., 2022), such as assuming the genders of people in stereotypically-male or stereotypically female occupations, even if the context unambiguously specifies otherwise  (Zhao et al., 2018; Rudinger et al., 2018; Parrish et al., 2021). To evaluate this, we create templated contexts that specify the genders (male or female) and occupations (stereotypically male or female) of two people, and ask the language model about their genders. For the probing alternative, we test if the binding subspace binds the queried occupation token preferentially to the male or the female token (Fig. 6 left). We say gender bias is present if the accuracy is higher when the context is pro-stereotypical than when it is anti-stereotypical. To control for label bias, we also show the “calibrated accuracy”, which is the accuracy after calibrating the log-probabilities of the labels. See Appendix H for details.
@@
Figure 6: Left: Anti-stereotypical example. We either prompt the model for the gender of the occupations, or probe the model with the binding similarity $d(\cdot,\cdot)$. Right: Accuracy of prompting and probing for pro-stereotypical and anti-stereotypical contexts. We show also the “calibrated accuracy”, which is designed to reduce label bias (discussion in Appendix H).
@@ 7 Conclusion
This work presents evidence for two hypotheses: first, that LMs internally construct symbolic models of input contexts; and second, when LMs are influenced by unfaithful tendencies, these internal models may remain faithful to the input context even if the outputs of the LMs are not.
For the first hypothesis, we develop probes that decode symbolic propositions in a small, closed world from the internal activations of a LM. Our work is primarily enabled by the discovery ofs the binding mechanism in LMs—we believe that this approach could be scaled to larger worlds with more complex semantics if more aspects of how LMs represent meaning are discovered, such as representations of role-filler binding (Smolensky, 1990) and state changes (Kim & Schuster, 2023).
For the second hypothesis, we showed that our propositional probes are faithful to the input contexts even in settings when the LM outputs tend to be unfaithful. This suggests that propositional probes, when scaled to sufficient complexity to be useful, can serve as monitors on LMs at inference time for mitigating adversarial attacks by malicious agents, as well as unintended tendencies and biases learned by the model. The latter could be more insidious—we often only discover surprising tendencies in models after we deploy them (Sharma et al., 2023; Pan et al., 2022; Roose, 2023).
@@ Acknowledgments
We thank Yossi Gandelsman, Shawn Im, Meena Jagadeesan, Xinyan Hu, Alexander Pan, and Lisa Dunlap for their helpful feedback on the manuscript. JF acknowledges support from the OpenAI Superalignment Fellowship. JS was supported by the National Science Foundation under Grants No. 2031899 and 1804794. In addition, we thank Open Philanthropy for its support of both JS and the Center for Human-Compatible AI.
@@ Appendix A Datasets
The name domain consists of 60 common English first names, which are all one-token wide in the llama 2 tokenizer. They are: The country domain consists of 16 countries, which are all one-token wide. They are: The food domain consists of 41 common foods, which are all two-tokens wide. They are: The occupation domain consists of 14 occupations, which are the subset of occupations used in the Winobias dataset (Zhao et al., 2018) (MIT license) that are one-token wide. They are: The para dataset is constructed by instructing GPT-3.5-turbo to rewrite the synth dataset into a story. The instructions used are: The trans dataset is constructed by instructing GPT-3.5-turbo to translate the para dataset into Spanish. The instructions used are:
@@ Appendix B Experimental details
Models   We use the huggingface implementation Wolf et al. (2019) as well as the TransformerLens library to run the Tulu-2-13B (Ivison et al., 2023) and Llama-2-13B-chat models (Touvron et al., 2023).
@@ Appendix C Details for the Hessian algorithm
The names and countries are random samples from the name and country domains. To reduce noise, we use 20 contexts, each constructed the same way. $F(x,y)$ itself is the average accuracy over these 20 contexts. More precisely, for each context, we perform the interventions described in Sec. 5.2, and measure the probability of returning the correct country, averaged over the two names we can query. We then average this across the 20 contexts. Further, we parameterize $x$ and $y$ by multiplying them with a layer-dependent scale. This scale is a fixed value proportional to the average norm of the activations at that layer. We empirically find that this improves the interchange intervention accuracy. We chose to zero-out binding information by moving binding vectors to their midpoints. We could also have chosen to change $E_{0},A_{0}$ to match $E_{1},A_{1}$, or vice versa. Empirically, mid-point works best of the three.
C.1 Binding subspace design choices
Some design choices in the binding subspace are informed by circuit-level analyses first conducted by Prakash et al. (2024). We first describe the high-level circuit they proposed, and then discuss why this informed the following choices:
  1. 1.
    Using a shared parameterization for $x$ and $y$ across layers for the Hessian algorithm (Sec 5.2)
  2. 2.
    Using a symmetric binding similarity metric (Eq. 1) based on ${\bm{U}}$ at layer $l=15$.
Figure 7: High-level circuit for binding mechanism. Resolving Bob’s country requires accessing binding information at two different layers: at the middle layers (around 15), the binding ID at the “Bob” token is retrieved, and at a later layer (around 30) the binding ID is used to look up the attribute that has the same binding ID as Bob, which in this case is “Peru”.
At a high level, there are two steps to resolving the query “Where does Bob live?” for the context “Alice lives in Laos. Bob lives in Peru.” (Fig. 7). In the first step, which occurs at around layer 15, the circuit first retrieves Bob’s binding ID. In the second step, Bob’s binding ID is used to look up attributes with the corresponding binding ID, in this case finding “Peru”.
However, we empirically find that the Peru’s binding vector (denoted lilac) and Bob’s binding vector (denoted green) are not quite the same vector. The two design choices have to factor this in.
We use a shared parameterization for $x$ and $y$ across all layers because for either “Bob” or “Peru”, the binding vector is only ever accessed either at around 15 or around 30, but never both. In principle, a more precise way of identifying the Hessian would be to identify the exact layers at which the binding vectors are accessed in “Bob” and “Peru”, and inject the perturbations $x$ and $y$ at those layers. However, we can get away with injecting the perturbations at all layers for both “Bob” and “Peru”, because the binding vectors are only accessed at layers 1̃5 and 3̃0 for the two tokens. This reduces the number of hyperparameters we have to tune. Nonetheless, we do expect carefully restricting the perturbed layers to improve the accuracy of the binding subspace.
We use a symmetric binding similarity metric because we hypothesize that the binding vectors for both “Bob” and “Peru” are the same at around layer 15, but evolve across the subsequent layers. This hypothesis explains the difference between the binding vectors of “Bob” and “Peru”, because the binding vector for “Bob” is extracted at around layer 15, but at around layer 30 for “Peru”. This implies that if we were to identify bound tokens using their layer 15 activations, we have to use the subspace corresponding the the layer 15 binding vectors, which according to the Hessian algorithm would be the left singular vectors ${\bm{U}}$.
Despite the circuit-motivated intuitions, the design choices are ultimately validated by the performance of propositional probes constructed from the resultant binding similarity metric.
@@ D.1 Distributed Alignment Search Baseline
In this section, we provide details of our implementation of Distributed Alignment Search. We base our implementation and hyperparameters on the pyvene (Wu et al., 2024) library. We use the Adam optimizer (Kingma & Ba, 2014), with learning rate 0.001, with a linear schedule over 5 epochs (with the first 0.1 steps as warmup), over a dataset of 128 samples, with batch size 8. We optimize over a subspace parametrized to be orthogonal using Householder reflections as implemented in pytorch (Paszke et al., 2019). This subspace is shared across all layers. The loss we use is the log probability of returning the desired attribute after performing the interchange intervention.
@@ Appendix F Grad-CAM attribution
In this section we describe the Grad-CAM style attribution we use to attribute information about domain values to context activations at specific layers and token positions. It is a general attribution technique that is originally invented for attributing information to pixels in the input space (Selvaraju et al., 2017), but has adapted for attributing information to internal activations (Olah et al., 2018).
@@ Appendix G Domain probes
We use Grad-CAM-style attribution to estimate which layers and which tokens carry the domain value information. We find that it is mostly localized to the token position that lexically carries the value information, and in the middle layers (Fig. 10 left). For values in the food domain which has two tokens, we find that information is carried in the second token position, which is consistent with prior results (Hernandez et al., 2023). We thus choose $l=20$ as the layer to probe from.
@@ Appendix H Gender bias evaluations
To ensure that the probing and prompting methods are detecting binding, and not relying on short cuts such as sentence order, for half of the contexts we swap the order of the last two sentences. We compare probing and prompting at predicting correctly the gender of an occcupation mentioned in the context. For prompting, we prompt the model with “The gender of the [occupation] is”, and take the gender with the higher log probability to be the model’s answer. For probing, we take the gender token with higher binding similarity to the occupation token to be the probe’s answer. To evaluate these two methods, we showed both accuracy and calibrated accuracy. Accuracy is the fraction of the time that asking for the gender of an occupation in the context returns the correct answer. Calibrated accuracy requires more explanation. Language models sometimes exhibit innocuous but systematic preferences when evaluated in a forced-choice setting. For example, it might encounter the “male” token a lot more frequently than the “female” token, and so it might output “male” over “female” regardless of what the context or even the bias in the occupation indicates. A common practice is to calibrate the log probabilities (Robinson et al., 2022). We do so by subtracting the mean log probabilities in paired responses. Specifically, let $v_{0},v_{1}\in\mathbb{R}^{2}$ be the log probabilities over the male and female tokens when queried with occupation 0 and occupation 1. The calibrated log probabilities for occupation 0 is $v_{0}-(v_{0}+v_{1})/2$, and that for occupation 1 is $v_{1}-(v_{0}+v_{1})/2$. After obtaining the calibrated log probabilities, we apply the same decision rule as before, i.e. we choose the gender with the higher calibrated log probability as the answer. The same procedure can be applied to the binding similarities to calibrate probing. However, we find that calibration does not significantly change the accuracy of either method.
Appendix I Quantitative position and order analyses
In this section, we perform a more thorough analysis of the effectiveness of propositional probes as we vary the position and order of the name and attribute tokens. Broadly speaking, the model could utilize several features to inform its representation of how information is bound: the model could represent binding as semantically conveyed by the context, or rely on spurious features such as position information (i.e. nearby tokens are bound) or order information (i.e. names appearing in an order are bound to attributes in the same order.)
Overall, we find that the binding subspace we extract is not sensitive to position, but is affected partially by order. However, we note that the model’s prompting performance becomes fragile when information is presented in unusual order, suggesting that the binding subspace we extract is influenced by order because the model’s own representations are in fact influenced by order. Further, we find that despite the partial susceptibility to order, the propositional probes still outperform prompting in the adversarial settings.
I.1 Position analysis
As suggested by anonymous reviewers, we construct datasets of varying lengths between the entity and attribute positions. Specifically, we collect 20 noun phrases such as “dedicated advocate”, 20 verb phrases such as “cultivates rare plants”, and compositionally create contexts of the form: “Alice, a dedicated advocate who cultivates rare plants, lives in Germany. Bob lives in France”. We call these contexts the long dataset. The medium dataset is similar, but does not contain verb phrases, e.g. “Alice, a dedicated advocate, lives in Germany. Bob lives in France.”, and the short dataset contains neither noun phrases nor verb phrases. As before, we sample random names and countries (as well as noun and verb phrases if applicable) to populate these templates, obtaining datasets of 512 input contexts.
Figure 11: Propositional probes do not degrade with different lengths between name and attribute tokens.
Fig. 11 shows that propositional probes do not degrade in long or medium datasets as compared to short. We additionally experimented with the version where both sentences in long have the long sentence structure, instead of just the first, and did not find a difference.
I.2 Order analysis
In this section, we construct templates in which the order of the names and entity tokens is varied. To reduce clutter, we introduce a notation that represent names as capital letters A or B, and their corresponding countries as numbers 1 or 2. Then, the order of a context could be written succinctly as strings such as “A1B2”. We also give each template a readable, informative name.
We list the templates and their examples below:
  1. 1.
    series (A1B2): “Alice lives in France. Bob lives in Germanay.”
  2. 2.
    cross (AB12): “Alice and Bob live in France and Germany respectively.”
  3. 3.
    reverse (A12B): “Alice lives in France. Germany is where Bob lives.”
  4. 4.
    coref (AB12): “Alice and Bob are friends. The former lives in France. The latter lives in Germany.”
  5. 5.
    nested (AB21): “Alice and Bob are friends. The latter lives in Germany. The former lives in France.”
Figure 12: Propositional probes do well in all data orderings except for nested.
We find that propositional probes are robust to all data orderings we studied except for nested (Fig. 12). On the nested dataset, probes have an EM accuracy of 44%. If the binding is done independently and at random, we expect the accuracy to be 25%. If the binding is done based on order, we expect the accuracy to be 0%. We thus interpret the 44% as saying that the the binding subspace captures both order information and the true semantic binding information.
We perform additional analyses to support this conclusion. First, we find that of the 56% where the probe outputs the wrong set of propositions, in 78% of these cases the probe assigns both countries to the same name. This suggests a continuous picture where the order information is in fact captured by our binding subspace, just that in some cases it is sufficiently strong enough to influence the decision boundary, and in others it is not. We find that a simple change where the propositions are constrained to bound to unique entities improves the accuracy to 76% (Fig. 13).
Figure 13: Constraining propositions to unique entities improves accuracy on both nested datasets (“matched” probes). A quirk in the language model makes the prompting strategy fail catastrophically on the second nested dataset.
Further, there is evidence that the model itself relies on shortcuts such as order information to capture binding. One evidence comes from an alternate nested dataset we constructed, nested_2 (AB21), that looks like: “Alice, unlike Bob who lives in Germany, lives in France.” We find that on this dataset, our prompting strategy fails catastrophically, whereas the probes have similar performance to the original nested dataset nested (Fig. 13). Our error analysis indicates that prompting fails because the model is confused about the order of the entities in the context, and tends to say that both the first name and the second name in the context are “Alice”. This suggests that in the nested order (AB21), the model’s internal representations of binding may be fragile and conflated with order.
Finally, even on the original nested dataset where prompting does not suffer from the catastrophic failure, and using the original propositional probes algorithm that does not enforce unique entities, we find that on the adversarial settings the probes still outperform prompting (Fig. 14).
Similarly, in the gender bias set up, we split the dataset into contexts in which binding is in the series order (AB12) and the nested order (AB21), and find that while probes perform worse in nested order than in series order, they still outperform prompting when the occupations are anti-stereotypical (Fig. 15).
Figure 14: In the adversarial prompt injection (P) and backdoor (FT) versions of the nested dataset, we find that probes still outperform prompting.
Figure 15: In the gender bias set up, the probe accuracy does degrade on the nested ordering (AB21) compared to the series ordering (AB12), but still outperforms prompting in both cases when examples are anti-stereotypical.
Appendix J Llama results
In this section, we show results for the Llama-2-13b-chat model, which is the instruction-tuned version of the base Llama-2-13b model. Our results are mostly similar.
Figure 16: (Llama) The accuracy of swapping binding information in name (attribute) activations by projecting into ${\bm{U}}_{(k)}$ (${\bm{V}}_{(k)}$) against $k$ in a context with 3 names and 3 attributes. We test the subspaces from the Hessian (blue), a random baseline (orange), and a skyline subspace obtained by estimating the subspace spanned by the first 3 binding vectors. We perform all 3 pairwise switches: 0-1 represents swapping the binding information of $E_{0}$ and $E_{1}$ ($A_{0}$ and $A_{1}$), and so on.
Figure 17: (Llama) Similarity between token activations under the binding similarity metric $d(\cdot,\cdot)$ for two-entity serial (left) and parallel (right) contexts.
(a) Llama domain probe accuracies
(b) Llama propositional probe accuracies
(a) Llama prop. probes in prompt injection (p) and dataset poisoning (ft) settings
(b) Llama prop. probes vs prompting for gender bias.